# On the universal validity of Case B recombination theory

### v1

@@ Abstract
keywords:
galaxies: starburst — galaxies: ISM — galaxies: dwarf1 Introduction
@@
Over the past decade, it has become clear that the galaxies we are able to observe in, or close to, the reionization epoch are characterized by young starbursts with strong nebular emission lines, compact sizes, and low dust content (e.g., Schaerer & de Barros, 2009; Finkelstein et al., 2012; Dunlop et al., 2013; Capak et al., 2015; Bouwens et al., 2017; De Barros et al., 2019; Endsley et al., 2020). Wide field optical surveys like the Sloan Digital Sky Survey (SDSS Blanton et al., 2017) have shown that galaxies can be identified at low- that have properties, where measurable, that are comparable to these reionization era systems. These analogs are characterized by compact morphologies and extreme nebular spectra (Cardamone et al., 2009), with high [O iii]/[O ii] line ratios (typically above 5, Izotov et al., 2017) and extreme Balmer line equivalent widths (EWs, over a few hundreds Å for H). Additionally, UV spectra of these objects show a high incidence of Ly in emission (very close to 100%), much higher than in normal star forming galaxies with similar stellar masses (e.g. Henry et al., 2015; Jaskot et al., 2019). These properties are indicative of low-metallicity (Z10% Z⊙, Cardamone et al., 2009; Amorín et al., 2010; Yang et al., 2017b; Senchyna et al., 2017), highly excited gas, ionized by young and hot starbursts, and with typically small amount of dust (e.g., Henry et al., 2018; Jaskot et al., 2017, 2019). Finally, over the past couple of years, direct measurements of their ionizing spectra (below rest–frame 912Å) proved that a substantial fraction of their ionizing radiation is able to escape the interstellar medium, with escape fraction up to 70%, (Izotov et al., 2016b; Izotov et al., 2016a; Schaerer et al., 2016; Verhamme et al., 2017; Izotov et al., 2018).
@@
Many of the physical properties derived from the emission line spectra of these galaxies are based on one fundamental assumption: that the Case B recombination limit is valid and can be used to interpret their spectra. Although this assumption is verified in most star-forming galaxies, there has been some evidence that the spectra of extreme emission line galaxies (EELGs) may deviate from Case B: 32 out of 41 of the EELGs in Yang et al. (2017b) have H/H ratio smaller than (H/H)CaseB. Similarly, Atek et al. (2009) find that a number of Ly emitters have comparably low Balmer line ratios. As part of an ongoing search aimed at pushing the selection of local analogs to lower masses and lower metallicities, described in detail in Mehta et al. (2020, in preparation), we discovered one object (hereafter, SXDF308) that shows an emission line spectrum that could dramatically change the physical interpretation of EELG spectra. The Balmer H/H line ratio measured in this object suggests that the universally accepted Case B limit is in fact not a good assumption for this galaxy. We argue in Section 4 that the physical conditions specific to this galaxy may in fact be more common than previously assumed.
@@ 2 Observations and data reduction
The spectroscopic data for SXDF308 were obtained in queue mode on Sept. 26, 2017 using the Hectospec Multifiber Spectrograph at the MMT 6.5m telescope in Arizona. Hectospec uses the secondary mirror, which incorporates atmospheric dispersion compensation prisms to correct for atmospheric dispersion between airmass of 1 and 2. We used the 270 mm-1 grating, blazed at 5200Å to provide a spectral coverage of 3650–9000Å (Fabricant et al., 2005), at a spectral resolution, for the used fiber diameter (15), of Å. Second-order blocking filters are not available to use with Hectospec. However, as we discuss below, the reddest line of interest ([S ii] Å) falls at 7200 Å, where the second order contamination from blue light is negligible (Gezari et al., 2012). The total integration time was 4,500 seconds, split in 5 exposures to remove cosmic ray hits. The seeing full-width-half-maximum during the observations was stable at 10, conditions were clear, and the average airmass of the observations was 1.4. A large number (103) of fibers were placed on empty sky111Empty sky regions were identified using SPLASH deep imaging regions, resulting in an accurate determination of the sky background, with the sky fiber-to-fiber variation of 0.2%.
@@ 2.1 Impact of observational effects
We estimate the contribution of chromatic seeing following the recent work of Herenz et al. (2017). These authors use Integral Field Spectroscopic (IFU) data to characterize the wavelength dependency of the seeing FWHM. Using their best fit linear fit (Table 2 in Herenz et al., 2017), we find that the seeing is 9% larger at 5200Å compared to 7000Å (the observed wavelengths of H and H respectively). Accordingly, given the 10 seeing during the observations, the energy enclosed by the 15 Hectospec fibers is 7% smaller at 5200Å compared to 7000Å. These calculations are valid for point sources. Given the uncertainty on the real galaxy profile, we do not apply this correction to the observed line ratio. We however caution the reader that the real H/H could be as low as 2.43, as a consequence of chromatic seeing.
@@ 5 Conclusions
SXDF308, together with the other examples among extreme emission line galaxies (Flury et al. 2020, in prep.) stand out and are identified as problematic because of their "unphysical" low Balmer H/H ratio, which in our explanation, is likely due to the particular geometry of the scattering gas. In general, though, galaxies for which radiative transfer effects of Balmer photons are important cannot be identified on the basis of this ratio alone, as, in the more general predictions, their values overlaps with the range expected for Case B plus extinction by dust. The addition of higher order Balmer lines and Paschen lines can help identifying these cases. More detailed modeling and additional observations are required to understand the impact of Balmer self-absorption to the interpretation of galaxies’ spectra.

### v2

@@ Abstract
1 Introduction
Over the past decade, it has become clear that the galaxies we are able to observe in, or close to, the reionization epoch are characterized by young starbursts with strong nebular emission lines, compact sizes, low dust content, and high ionization parameters (e.g., Schaerer & de Barros, 2009; Finkelstein et al., 2012; Dunlop et al., 2013; Capak et al., 2015; Bouwens et al., 2017; De Barros et al., 2019; Endsley et al., 2020). These results have now been confirmed with extensive spectroscopic campaigns conducted with the James Webb Space Telescope, which push observations of rest-frame optical lines out to redshifts (e.g., Topping et al., 2024; Pirzkal et al., 2023).
@@ 1 Introduction
Wide field optical surveys like the Sloan Digital Sky Survey (SDSS Blanton et al., 2017) have shown that galaxies can be identified at low- that have properties, comparable to these reionization era systems (Schaerer et al., 2022). These analogs are characterized by compact morphologies and extreme nebular spectra, with high [O iii]/[O ii] line ratios (typically above 5, Izotov et al., 2017) and extreme Balmer line equivalent widths (EWs, over several hundreds Å for H, and [O iii]). Additionally, UV spectra of these objects show a high incidence of Ly in emission (very close to 100%), much higher than in normal star forming galaxies with similar stellar masses (e.g. Henry et al., 2015; Jaskot et al., 2019). These properties are indicative of low-metallicity (Z10% Z⊙, Cardamone et al., 2009; Amorín et al., 2010; Yang et al., 2017b; Senchyna et al., 2017), highly excited gas, ionized by young and hot starbursts, and with typically small amount of dust (e.g., Henry et al., 2018; Jaskot et al., 2017, 2019). Finally, direct measurements of their ionizing spectra (below rest–frame 912Å) proved that a substantial fraction of their ionizing radiation is able to escape the interstellar medium, with escape fraction up to 70%, (Izotov et al., 2016b; Izotov et al., 2016a; Schaerer et al., 2016; Verhamme et al., 2017; Izotov et al., 2018; Flury et al., 2022b, a).
@@
Many of the physical properties derived from the emission line spectra of these galaxies are based on one fundamental assumption: that the Case B recombination limit is valid and can be used to interpret their spectra. Although this assumption is verified in the vast majority of star-forming galaxies, there has been some evidence that the spectra of extreme emission line galaxies (EELGs) may deviate from Case B: 32 out of 41 of the EELGs in Yang et al. (2017b) have H/H ratio smaller than (H/H)CaseB. Similarly, Atek et al. (2009) find that a number of Ly emitters have comparably low Balmer line ratios. Early results based on JWST spectra show similar trends: Pirzkal et al. (2023) find a significant number of sources in the NGDEEP-NISS survey which appear to violate Case B recombination. This result seems to hold at even higher redshifts, where examples of H/H line ratios smaller than Case B are found at by, e.g., Cameron et al. (2023) and Topping et al. (2024). As part of an ongoing search aimed at pushing the selection of local analogs to lower masses and lower metallicities, described in Lin et al. (2022), we discovered one object (hereafter, SXDF308) that shows an emission line spectrum that could dramatically change the physical interpretation of EELG spectra. The Balmer H/H line ratio measured in this object suggests that the universally accepted Case B limit is in fact not a good assumption for this galaxy. We argue in Section 4 that the physical conditions specific to this galaxy may in fact be more common than previously assumed.
@@ 2 Observations and data reduction
[…]
@@ 2.1 Impact of observational effects
[…]
@@ 5 Conclusions
SXDF308, together with the other examples among extreme emission line galaxies stand out and are identified as problematic because of their "unphysical" low Balmer H/H ratio, which in our explanation, is likely due to the particular geometry of the scattering gas. In general, though, galaxies for which radiative transfer effects of Balmer photons are important cannot be identified on the basis of this ratio alone, as, in the more general predictions, their values overlaps with the range expected for Case B plus extinction by dust. The addition of higher order Balmer lines and Paschen lines can help identifying these cases. More detailed modeling and additional observations are required to understand the impact of Balmer self-absorption to the interpretation of galaxies’ spectra.